# Swarm Intelligence in Geo-Localization: A Multi-Agent Large Vision-Language Model Collaborative Framework

### v2

@@ Abstract.
Visual geo-localization demands in-depth knowledge and advanced reasoning skills to associate images with real-world geographic locations precisely. In general, traditional methods based on data-matching are hindered by the impracticality of storing adequate visual records of global landmarks. Recently, Large Vision-Language Models (LVLMs) have demonstrated the capability of geo-localization through Visual Question Answering (VQA), enabling a solution that does not require external geo-tagged image records. However, the performance of a single LVLM is still limited by its intrinsic knowledge and reasoning capabilities. To address these challenges, we introduce smileGeo, a novel visual geo-localization framework that leverages multiple Internet-enabled LVLM agents operating within an agent-based architecture. By facilitating inter-agent communication, smileGeo integrates the inherent knowledge of these agents with additional retrieved information, enhancing the ability to effectively localize images. Additionally, our framework employs a dynamic learning strategy that optimizes communication among agents, minimizing redundant interactions and improving overall system efficiency. To validate the effectiveness of the proposed framework, we conducted experiments on three different datasets, and the results show that our approach significantly outperforms current state-of-the-art methods. The source code is available at https://anonymous.4open.science/r/ViusalGeoLocalization-F8F5.
@@ 1. Introduction
Visual geo-localization, referred to the task of estimating geographical identification for a given image, is vital in various fields such as analyzing historical human mobility patterns (Huang and Carley, 2019; Luo et al., 2011; Arandjelovic et al., 2018; Zaffar et al., 2021; Torii et al., 2018) and providing location-aware, city-level attraction recommendations (Chen et al., 2017a, 2018, b; Garg et al., 2022; Hausler et al., 2019; Khaliq et al., 2020). In general, accurately geo-localizing images without relying on localization-based metadata (e.g., GPS tags) is a complex task that demands extensive geospatial knowledge and advanced reasoning capabilities. Traditional methods (ElQadi et al., 2020; Campbell and Wheeler, 2006; Deng et al., 2020; Zhang et al., 2018) typically formulate it as an image retrieval problem where to geo-localize the given image by retrieving similar images with known geographical locations. Thus, their effectiveness is limited by the scope and quality of the geo-tagged image records.
@@
  • We conducted experiments on two open-source datasets. To address the issue of numerous images in these datasets that could not be localized, we also constructed a new private dataset for further evaluation. Extensive experimentation demonstrates that smileGeo achieves competitive performance compared to state-of-the-art methods.
@@ 4.1. Experiment Setup
Datasets. In this paper, we mainly evaluate the proposed geo-localization framework, smileGeo, on the two open-source datasets: IM2GPS3K [^1] http://www.mediafire.com/file/7ht7sn78q27o9we/im2gps3ktest.zip and YFCC4K [^2] http://www.mediafire.com/file/3og8y3o6c9de3ye/yfcc4k.zip:
The IM2GPS3K dataset is a widely used benchmark for visual geo-localization. It consists of 3,000 images from the Flickr photo-sharing platform, which are tagged with precise geographical coordinates. The dataset covers a diverse range of locations globally, including urban, rural, and natural environments, making it an ideal testbed for assessing the generalization capabilities of geo-localization models. The images in IM2GPS3K are drawn from various categories such as landscapes, cityscapes, and landmarks, providing a challenging and diverse set of visual cues that the models must recognize to predict the geographical location accurately.
The YFCC4K dataset is a subset of the larger Yahoo Flickr Creative Commons 100 Million (YFCC100M) dataset, which contains over 100 million Flickr images. The YFCC4K subset includes 4,000 images with corresponding geographical tags, carefully selected to represent a broad geographic distribution and visual diversity. Like IM2GPS3K, YFCC4K includes a variety of image types, such as natural landscapes, urban settings, and iconic landmarks. This diversity allows for comprehensive evaluation of geo-localization models across different environments and scales, from global-level to city-level localization tasks.
@@
In addition, we newly construct a geo-localization dataset named GeoGlobe. It contains a variety of man-made landmarks or natural attractions from nearly 150 countries with different cultural and regional styles. The diversity and richness of GeoGlobe allow us to evaluate the performance of different models more accurately. More details can be found in Appendix B.
@@ 4.3. Ablation Study
To demonstrate the benefits of allowing every LVLM agent to retrieve information from the Internet for enhancing the knowledge of the model, as well as the potential for multiple LVLMs to use the retrieved data for further reasoning, we designed an experiment as shown in Figure LABEL:fig:retrieval. Figure LABEL:fig:retrieval_a illustrates the results without any retrieval, while Figure LABEL:fig:retrieval_b displays the results when models incorporate the retrieved information. We compare the proposed model with two types of methods: image similarity-based retrieval approaches and advanced closed-source large models. The results indicate that image similarity-based methods, such as Translocator and GeoCLIP, struggle to effectively leverage Internet-retrieved information, leading to minimal changes in accuracy. In contrast, closed-source large models show moderate performance improvements when aided by additional retrieval data. Notably, our method consistently outperforms all baselines across datasets, with a significant improvement of over 30% in the GeoGlobe (natural) dataset for smileGeo. This demonstrates that our proposed framework, with its robust reasoning capabilities, can fully exploit retrieved information to enhance reasoning and model accuracy.
@@ 4.4. Efficiency Study
Table 4.4 reports the average and median of response times ($RT_{\text{Avg}}$, $RT_{\text{Med}}$), as well as the model accuracy ($Acc$). In general, as the maximum allowable discussion round $Z$ increases, the response time of smileGeo also increases. However, thanks to the implementation of our dynamic learning model with an early stopping strategy, this growth remains nearly linear, rather than exponential.
Then, the comparison of $RT_{\text{Avg}}$ and $RT_{\text{Med}}$ shows that only a small fraction of image queries require multiple discussion rounds. These outliers, with significantly longer response times, skew the average downward. From the median response time, we observe that at least 50% of the test samples receive a response within 2 seconds, an acceptable performance level for most applications.
@@ 4.5. Parameter Analysis
Hyperparameter $K$ $R$. There are two hyperparameters, $K$ and $R$, that need to be pre-defined in the proposed framework: $K$ is the number of agents (answer agents) that respond in each round of discussion, and $R$ is the number of agents (review agents) used to review answers from answer agents. Therefore, we conduct experiments under different combinations of $K\in[1,8]$ and $R\in[1,8]$, as shown in Figure 3. The results indicate that optimal performance can be achieved with relatively small values of $K$ or $R$. However, the computational cost, measured in tokens, increases exponentially with higher values of $K$ and $R$. To balance both the efficiency and the accuracy of smileGeo, for the experiments presented in this paper, we set both $K$ and $R$ equal to 2.
@@ 5. Conclusion
Looking ahead, we aim to expand the capabilities of smileGeo to incorporate more powerful external tools beyond just web searching. Additionally, we plan to explore extending its application to complex scenarios, such as high-precision global positioning and navigation for robots, laying the cornerstone for exploring LVLM agent collaboration to handle different complex open-world tasks efficiently. More discussions are detailed in Appendix E.
@@ Appendix D Case Study
Case 1: In Figure 5, we illustrate the application of smileGeo in a visual geo-localization task. For this demonstration, we randomly select an image from the test dataset and employ five distinct LVLMs: LLaVA, GPT-4o mini, Claude-3, Gemini, and Qwen. The agent selection model selects two answer agents, as depicted in the top part of the figure. Subsequently, stages 1 through 3 detail the process of generating the accurate geo-location. Initially, only one answer agent provided the correct response. However, after several rounds of discussion, the agent that initially responded incorrectly revised the confidence level of its answer. During the final internal discussion, this agent aligned its response with the correct answer. This outcome validates the efficacy of our proposed framework, demonstrating its ability to integrate the knowledge and reasoning capabilities of different agents to enhance the overall performance of the proposed LVLM agent framework.

### v3

@@ Abstract.
Visual geo-localization demands in-depth knowledge and advanced reasoning skills to associate images with precise real-world geographic locations. Existing image database retrieval methods are limited by the impracticality of storing sufficient visual records of global landmarks. Recently, Large Vision-Language Models (LVLMs) have demonstrated the capability of geo-localization through Visual Question Answering (VQA), enabling a solution that does not require external geo-tagged image records. However, the performance of a single LVLM is still limited by its intrinsic knowledge and reasoning capabilities. To address these challenges, we introduce smileGeo, a novel visual geo-localization framework that leverages multiple Internet-enabled LVLM agents operating within an agent-based architecture. By facilitating inter-agent communication, smileGeo integrates the inherent knowledge of these agents with additional retrieved information, enhancing the ability to effectively localize images. Furthermore, our framework incorporates a dynamic learning strategy that optimizes agent communication, reducing redundant interactions and enhancing overall system efficiency. To validate the effectiveness of the proposed framework, we conducted experiments on three different datasets, and the results show that our approach significantly outperforms current state-of-the-art methods. The source code is available at https://anonymous.4open.science/r/ViusalGeoLocalization-F8F5.
@@ 1. Introduction
Visual geo-localization, referred to the task of estimating geographical identification for a given image, is vital in various fields, such as analyzing historical human mobility patterns (Huang and Carley, 2019; Luo et al., 2011; Arandjelovic et al., 2018; Zaffar et al., 2021; Torii et al., 2018) and providing location-aware, city-level attraction recommendations (Chen et al., 2017a, 2018, b; Garg et al., 2022; Hausler et al., 2019; Khaliq et al., 2020). In general, accurately geo-localizing images without relying on localization-based metadata (e.g., GPS tags) is a complex task that demands extensive geospatial knowledge and advanced reasoning capabilities. Traditional methods (ElQadi et al., 2020; Campbell and Wheeler, 2006; Deng et al., 2020; Zhang et al., 2018) typically formulate it as an image retrieval problem where to geo-localize the given image by retrieving similar images with known geographical locations. Thus, their effectiveness is limited by the scope and quality of the geo-tagged image records.
@@
  • We conducted experiments on two open-source datasets. To address the issue of numerous images (food, furniture, etc. ) in these datasets that could not be localized, we also constructed a new dataset for further evaluation. Extensive experimentation demonstrates that smileGeo achieves competitive performance compared to state-of-the-art methods.
@@ 4.1. Experiment Setup
Datasets. In this paper, we first evaluate the proposed geo-localization framework, smileGeo, on the two open-source datasets: IM2GPS3K [^1] http://www.mediafire.com/file/7ht7sn78q27o9we/im2gps3ktest.zip and YFCC4K [^2] http://www.mediafire.com/file/3og8y3o6c9de3ye/yfcc4k.zip:
The IM2GPS3K dataset is a widely used benchmark for visual geo-localization. It consists of 3,000 images from the Flickr photo-sharing platform, which are tagged with precise geographical coordinates. The dataset covers a diverse range of locations globally, including urban, rural, and natural environments, making it an ideal testbed for assessing the generalization capabilities of geo-localization models. The images in IM2GPS3K are drawn from various categories, such as landscapes, cityscapes, and landmarks, providing a challenging and diverse set of visual cues that the models must recognize to predict the geographical location accurately.
The YFCC4K dataset is a subset of the larger Yahoo Flickr Creative Commons 100 Million (YFCC100M) dataset, which contains over 100 million Flickr images. The YFCC4K subset includes 4,536 images with corresponding geographical tags, selected to represent a broad geographic distribution and visual diversity. Like IM2GPS3K, YFCC4K includes a variety of image types, such as natural landscapes, urban settings, and iconic landmarks. This diversity allows for comprehensive evaluation of geo-localization models across different environments and scales, from global-level to city-level localization tasks.
@@
In addition, We have newly constructed a geo-localization dataset named GeoGlobe. It contains a variety of man-made landmarks or natural attractions from nearly 150 countries with different cultural and regional styles. The diversity and richness of GeoGlobe allow us to evaluate the performance of different models more accurately. More details can be found in Appendix B.
@@ 4.3. Ablation Study
To demonstrate the benefits of allowing every LVLM agent to retrieve information from the Internet for enhancing the knowledge of the model, as well as the potential for multiple LVLMs to use the retrieved data for further reasoning, we designed an experiment as shown in Figure LABEL:fig:retrieval. Figure LABEL:fig:retrieval_a illustrates the results without any retrieval, while Figure LABEL:fig:retrieval_b displays the results when models incorporate the retrieved information. We compare the proposed model with two types of methods: image similarity-based retrieval approaches and advanced closed-source large models. The results indicate that image similarity-based methods, such as Translocator and GeoCLIP, struggle to effectively leverage Internet-retrieved information, leading to minimal changes in accuracy. In contrast, closed-source large models show moderate performance improvements when aided by additional retrieval data. Notably, our method consistently outperforms all baselines across datasets, with a significant improvement of over 20% in the GeoGlobe (natural) dataset for smileGeo. This demonstrates that our proposed framework, with its robust reasoning capabilities, can fully exploit retrieved information to enhance reasoning and model accuracy.
@@ 4.4. Efficiency Study
Table 4.4 reports the average and median of response times ($RT_{\text{Avg}}$, $RT_{\text{Med}}$), as well as the model accuracy ($Acc$). In general, as the maximum allowable discussion round $Z$ increases, the response time of smileGeo also increases. However, thanks to implementing our dynamic learning model with an early stopping strategy, this growth remains nearly linear rather than exponential.
Then, the comparison of $RT_{\text{Avg}}$ and $RT_{\text{Med}}$ shows that only a small fraction of image queries require multiple discussion rounds. These outliers, with significantly longer response times, skew the average upwards. From the median response time, we observe that at least 50% of the test samples receive a response within 2 seconds, an acceptable performance level for most applications.
@@ 4.5. Parameter Analysis
Hyperparameter $K$ $R$. There are two hyperparameters, $K$ and $R$, that need to be pre-defined in the proposed framework: $K$ is the number of agents (answer agents) that respond in each round of discussion, and $R$ is the number of agents (review agents) used to review answers from answer agents. Therefore, we conduct experiments under different combinations of $K\in[1,8]$ and $R\in[1,8]$, as shown in Figure 3. The results indicate that optimal performance can be achieved with relatively small values of $K$ or $R$. However, the computational cost, measured in tokens, increases exponentially with higher values of $K$ and $R$. To ensure the accuracy of the smileGeo while reducing the calculation cost as much as possible, we set both $K$ and $R$ equal to 2 in this paper.
@@ 5. Conclusion
Looking ahead, we aim to expand the capabilities of smileGeo to incorporate more powerful external tools beyond just web searching. Additionally, we plan to explore extending its application to more complex scenarios, such as high-precision global positioning, laying the cornerstone for exploring LVLM agent collaboration to handle different complex open-world tasks efficiently. More discussions are detailed in Appendix E.
@@ Appendix D Case Study
Case 1: In Figure 5, we illustrate the application of smileGeo in a visual geo-localization task. For this demonstration, we randomly select an image from the test dataset and employ five distinct LVLMs: LLaVA, GPT-4o-mini, Claude-3-opus, Gemini-1.5-pro, and Qwen2. The agent selection model selects two answer agents, as depicted in the top part of the figure. Subsequently, stages 1 through 3 detail the process of generating the accurate geo-location. Initially, only one answer agent provided the correct response. However, after several rounds of discussion, the agent that initially responded incorrectly revised the confidence level of its answer. During the final internal discussion, this agent aligned its response with the correct answer. This outcome validates the efficacy of our proposed framework, demonstrating its ability to integrate the knowledge and reasoning capabilities of different agents to enhance the overall performance of the proposed LVLM agent framework.